# CELA: Cost-Efficient Language Model Alignment for CTR Prediction

### v2

@@ 4.1.2. ID-Based Model Pre-training
4.2. ROMA: Recommendation-Oriented Modal Alignment
Despite the profound domain understanding that the PLM achieves through pretraining, a significant gap persists between their output space and that of collaborative filtering. This discrepancy arises because the capabilities of the PLM extend well beyond what recommendation tasks require. Without targeted training and constraints, the outputs of the PLM may not align with the needs of recommender systems. Consequently, it is crucial to refine and orient the PLM’s outputs towards recommendation tasks, a process referred to as alignment. This essential alignment, which aims to provide a representation of the text feature that can be seamlessly integrated with other features for the next stage, is illustrated in the central portion of Figure 2. The substantial number of parameters in the PLM and the extensive volume of interactions in the dataset render interaction-level alignment inefficient (Li et al., 2023; Liu et al., 2022; Geng et al., 2022). Consequently, we propose item-level alignment as a solution.
First, the textual and non-textual features on the item side are encoded into their respective representations. For item $i$, its non-textual features $\boldsymbol{x}^{i,n}$, are encoded utilizing the embedding table from the ID-based model developed in the first stage. Conversely, the textual feature $\boldsymbol{x}^{i,t}$ is processed through the PLM, followed by a transformation via a projection head. This head is to adjust the dimensionality of the output from the PLM to match that of the non-textual features. This encoding process can be formulated as:
@@ 4.2. ROMA: Recommendation-Oriented Modal Alignment
where $\boldsymbol{e}^{n}_{i}$ signifies the encoded representation of item $i$’s non-textual features, and $\boldsymbol{e}^{t}_{i}$ represents the encoded textual feature. The terms $\mathbf{W}$ and $\boldsymbol{b}$ are the parameters of the projection head, comprising the matrix and bias, respectively.
Contrastive learning is then utilized to align the representations of textual and non-textual features. Within a given batch, representations of these features from the same item are considered positive pairs, while unmatched representations are treated as negative pairs. Given that both modalities depict the same item, there is an inherent overlap of insights between its positive pair. By incentivizing the model to differentiate between matching and mismatched pairs, contrastive learning facilitates the establishment of correspondence or mapping functions across modalities. The alignment begins with
(10) $\small\mathcal{L}_{t,n}=-\frac{1}{\left|\mathcal{I}\right|}\sum_{i\in\mathcal{I}}\log\frac{\exp\left(sim_{2}\left(\boldsymbol{e}^{t}_{i},\boldsymbol{e}^{n}_{i}\right)/\tau_{2}\right)}{\sum_{k\in\mathcal{I}}\exp\left(sim_{2}\left(\boldsymbol{e}^{t}_{i},\boldsymbol{e}^{n}_{k}\right)/\tau_{2}\right)},$
where $sim_{2}(\cdot)$ denotes the inner product similarity and $\tau_{2}$ denotes another temperature.
@@
(11) $\small\mathcal{L}_{n,t}=-\frac{1}{\left|\mathcal{I}\right|}\sum_{i\in\mathcal{I}}\log\frac{\exp\left(sim_{2}\left(\boldsymbol{e}^{n}_{i},\boldsymbol{e}^{t}_{i}\right)/\tau_{2}\right)}{\sum_{k\in\mathcal{I}}\exp\left(sim_{2}\left(\boldsymbol{e}^{n}_{i},\boldsymbol{e}^{t}_{k}\right)/\tau_{2}\right)}.$
The overall modal alignment loss combines these two components as:
@@
Training is exclusively confined to the PLM and projection head, with the embedding table remaining non-trainable to prevent disruption of the collaborative filtering space. It is important to emphasize that, the item-level alignment significantly reduces the computational resources and time required for training.
@@ 4.3. MF2: Multi-Modal Feature Fusion
To mitigate suboptimal performance caused by purely text-based recommendations, we preserve the collaborative signals captured by the ID-based model while integrating the PLM as a text encoder. This PLM is tailored to the domain of the textual feature and capable of aligning its outputs to the collaborative space. Serving solely as a text encoder, the PLM prevents scalability issues from increased features, and allows for a detailed description of user behaviors.
To reduce training time and online latency, we create an item text embedding table. Text for each item is processed through the aligned PLM to obtain the embedding, stored for quick access by item identifiers, eliminating the need for real-time PLM processing.
Then, as depicted on the right side of Figure 2, the recommender system is enhanced by integrating the textual feature alongside non-textual features. This textual feature, indexed from the text embedding table using the item’s identifier, is processed using a MLP to ensure dimensional consistency. Within the feature interactions layer, textual and non-textual features are treated uniformly. Such uniformity across all modalities prevents substantial modifications to the original network, thus facilitating its application to any network architecture. This process can be formulated as:
@@
where $\mathbf{E}^{i,t}$ denotes the text embedding table, $\boldsymbol{x}^{i,id}$ denotes the target item’s identifier, $\boldsymbol{x}^{h,id}$ denotes all the items’ identifiers in user behaviors, $FI(\cdot)$ denotes the feature interaction layer and $FC(\cdot)$ denotes the fully connected layer.
@@ 4.4. Alternate Training
The last two stages, as indicated by the green dotted line in Figure 2, are designed to be executed in an alternate fashion. Following the completion of the embedding table training in the MF2 stage, we can revert to the ROMA stage. Here, the item text representation is realigned to the newly trained embedding table. Subsequently, this realigned language model is utilized to generate a new text embedding table. This updated table is then employed in the subsequent iteration of the MF2 stage for further training. This alternate approach allows for iterative refinement and enhancement of both the alignment and fusion processes, thereby progressively improving the overall performance of the model.
@@ 5.1.1. Datasets
We conduct offline experiments on three datasets that include textual features. The statistics for these datasets are provided in Table 1.
[…]
MovieLens Dataset  [^2] https://grouplens.org/datasets/movielens/latest/ collectes the rating activity from MovieLens  [^3] http://movielens.org/. Due to the necessity to extract movie summaries as the textual feature from provided hyperlinks, we sample 10 million interactions through the random sampling of users. The task is to predict whether a user will like the next item, with a preference indicated by a rating of 3 or higher, based on a sequence of up to 20 user behaviors.
AppGallery Dataset comprises application exposure and download records sampled from Huawei AppGallery over eight days. Data from the first seven days constituted the training set, while data from the subsequent day served as the test set. The task is to predict whether the user will download the next application, with application description as the textual feature and a sequence of up to 14 user behaviors.
Table 1. Statistics of datasets for offline experiments.
Dataset #Users #Items #Features #Samples
Amazon 300,562 100,682 6 4,494,230
MovieLens 92,911 26,084 4 10,146,329
AppGallery 11,056,819 29,759 8 16,398,193
@@ 5.1.2. Baselines
Our method is benchmarked against contemporary state-of-the-art models, encompassing both ID-based and text-enhanced models. Classic ID-based models, such as AutoInt (Song et al., 2019), DCN (Wang et al., 2017), DCNv2 (Wang et al., 2021), FiBiNET (Huang et al., 2019), PNN (Qu et al., 2016) , WideDeep (Cheng et al., 2016), DeepFM (Guo et al., 2017), and xDeepFM (Lian et al., 2018), focus on modeling high-order feature interactions. In contrast, models like DIN (Zhou et al., 2018) and DIEN (Zhou et al., 2019) specialize in fine-grained user behavior modeling. Text-enhanced models, including P5 (Geng et al., 2022) and PTab (Liu et al., 2022), aim to describe data in natural language to elicit responses from language models. CTR-BERT (Muhamed et al., 2021) integrates traditional and textual features for fusion in a MLP, UniSRec (Hou et al., 2022) utilizes a frozen BERT (Devlin et al., 2018) as its text encoder, and CTRL (Li et al., 2023) aligns text and tabular data representations to infuse the ID-based model with semantic information.
@@ 5.1.3. Evaluation Metrics
In this research, Area Under the Curve (AUC) and logarithmic loss (Logloss) are adopted as evaluation metrics, both widely acknowledged in the assessment of CTR models (Song et al., 2019; Wang et al., 2021; Qu et al., 2016; Guo et al., 2017). AUC evaluates the model’s capacity to accurately rank a positive instance above a negative one, thereby indicating its ranking efficacy. Logloss, on the other hand, quantifies the degree of correspondence between the predicted and actual CTR, offering a metric for the model’s predictive accuracy.
5.1.4. Implementation Details
In our offline experiments, models are implemented using RecStudio (Lian et al., 2023) and Transformers (Wolf et al., 2020). We integrate RoBERTa (Liu et al., 2019) to enhance the optimal ID-based model. AdamW (Loshchilov and Hutter, 2017) optimizer facilitates all training stages. For the DAP stage, $\tau_{1}$ ranges from [0.001 to 0.1] and $\alpha$ from [0.1 to 1.5], with a batch size of 32. In the ROMA stage, $\tau_{2}$ is 1, batch size is 128, the language model’s learning rate is 5e-5, and projection units are [128, $f_{i}d$], with $f_{i}$ as item-side feature count and $d$ set at 16. The MF2 stage uses a batch size of 1024 and standardized hidden units [128, 64, 64] across ID-based models. ID-based models that do not focus on user behavior modeling aggregate user behavior sequences using mean pooling.
@@ 5.2. Overall Performance (RQ1)
  • Within ID-based models, DIN and DIEN excel on Amazon, because their fine-grained modeling of user behaviors caters to the pronounced reliance on user behaviors. Conversely, for Movielens and AppGallery, the dependency on user behaviors appears less critical, resulting in no marked advantage for DIN and DIEN.
  • CTRL ranks second to CELA in most cases due to its integration of the semantics into the ID-based model, distinguishing it from traditional ID-based models. Unlike other text-enhanced models, CTRL retains collaborative signals, enhancing its performance.
  • Among text-enhanced models, P5 and PTab perform worse than most ID-based models. This observation is consistent with results from (Li et al., 2023), corroborating that recommender systems based purely on text generally underperform compared to ID-based models.
  • CELA demonstrates superior performance across all datasets, thereby validating three core assertions: firstly, the utilization of the PLM extracts intrinsic knowledge within the textual feature; secondly, employing the PLM solely as a textual feature encoder prevents information loss caused by an increase in features and enables the detailed delineation of items; thirdly, the integration of collaborative signals significantly enhances model efficacy. This last point is further supported by the superior outcomes observed in the CTRL and UniSRec compared to other baselines.
Table 2. Performance comparison with baselines. The best and the second best results are in bold and underlined, respectively. denotes the statistical significance of the performance difference with respect to the second-best result at the 0.05 level.
@@ 5.3. Efficiency Measurement (RQ2)
To explore the practicality of CELA in industrial scenarios, we measure its training duration for one epoch and inference latency for a single batch on the industrial dataset AppGallery, compared to baselines. The training duration is calculated for the most time-consuming stage. Results presented in Table 3 reveals that:
  • Within text-enhanced models, the training expenses of P5, PTab, CTR-BERT, and CTRL are notably high, attributable to their complexity, which is implied by the number of parameters. the inference overhead of P5, PTab, and CTR-BERT is also very high for the same reason. Although reducing the update frequency can reduce training costs, employing an outdated language model adversely affects recommendation performance. Therefore, it is not practical to apply them in real-world scenarios.
  • UniSRec, by not fine-tuning the language model, maintains training and inference costs comparable to those of DIN. However, using a frozen language model limits its efficacy in recommendation tasks. This limitation stems from the insufficient alignment of the PLM’s knowledge with the specific requirements of the recommendation domain, resulting in suboptimal integration of textual and non-textual features.
  • The ROMA stage of CELA necessitates fine-tuning of the language model, but this tuning is conducted at the item level rather than the interaction level. Typically, the number of items is significantly lower than that of interactions, leading to minimal training costs, particularly in contexts with a limited number of items. The training and inference costs of CELA’s MF2 are on par with those of DIN, making CELA industry-friendly.
Table 3. Comparison of training and inference efficiency between different models on AppGallery. Params is the number of trainable parameters, the training time is calculated for the whole process and the inference time is calculated for one batch.
Model Params (M) Inference (ms) Total (h)
@@
CELA DAP (102.88+0.58) 103.46 / (1.07+1.76) 2.83
CELA ROMA 102.30 / 0.30
@@
CELA / / 5.42
@@
To assess CELA’s efficacy in cold-start scenarios, the test dataset is segmented into groups based on item popularity, with the AUC evaluated for each group. As illustrated in Figure 3, CELA consistently outperforms the baseline across all cold-start conditions. This superior performance is attributed to the incorporation of a PLM that leverages embedded external knowledge for text comprehension. Notably, when the popularity falls below 30, the improvements of CELA become particularly pronounced. This is because, under such conditions, ID-based models struggle to learn effective embeddings owing to their dependence on historical interactions.
@@ 5.5. Ablation Study (RQ4)
To explore the impact of different training processes and components on the overall performance of CELA, we conduct an ablation study with designed several variants:
  • DIN + RoBERTa: Incorporate RoBERTa as the text encoder and directly undergo the MF2 stage with DIN. It is equivalent to our CELA without the DAP and ROMA stages.
@@
  • CELA w/o ROMA: The ROMA stage is removed from CELA.
  • DAP w/o SimCSE: The SimCSE loss is removed in DAP.
  • DAP w/o MLM: The MLM loss is removed in DAP.
  • ROMA w/o PLM: A text embedding table encoded by the PLM replaces the PLM in ROMA.
  • ROMA w/ cos: Utilizing cosine similarity in ROMA.
  • MF2 w/o text: The text embedding table in MF2 is replaced by DAP’s embedding table from the ID-based model.
Analysis of Table 4 reveals the following insights:
  • DIN + RoBERTa yields fewer performance improvements compared to CELA, suggesting that the mere inclusion of textual feature is insufficient for effective information extraction and utilization in recommendation tasks. The extent of the performance impact varies with the degree of overlap between the PLM’s corpus and the textual features in the dataset.
  • Removing the DAP stage from CELA leads to a reduction in performance, as it deprives the PLM of the opportunity to comprehend domain-specific knowledge, and the anisotropy issue compromises the PLM’s output expressiveness. Nevertheless, the subsequent ROMA stage aligns the PLM output with the recommendation domain, enabling the recommender system to understand PLM outputs. Although the PLM’s knowledge is not perfectly tailored for the domain, it results in performance that still surpasses that of DIN.
  • The removal of the ROMA stage from CELA also leads to diminished performance. While the DAP stage augments the PLM’s capability to comprehend and express the textual feature, a discrepancy between the PLM output space and the collaborative space hinders the recommender system’s understanding of PLM outputs for CTR prediction.
  • Removing the SimCSE loss from DAP leads to a performance decline, as it overlooks the anisotropy issue present in PLMs, thereby diminishing expressiveness. Incorporating SimCSE loss is advisable to enhance the quality of text representations.
  • Removing the MLM loss from DAP results in diminished performance. This is because MLM, serving as a widely adopted pretraining loss, facilitates the swift adaptation of language models to specialized domains. DAP w/o MLM performs even worse than CELA w/o DAP on Amazon dataset, suggesting that using SimCSE loss only in pretraining is not ideal due to the inconsistency with the objective of PLMs’ training. Originally, SimCSE loss was proposed as an auxiliary objective in pretraining, aiming to regularize pre-trained embeddings’ anisotropic space to be more uniform.
  • Removing the PLM significantly reduces performance due to less semantic information preservation when using ROMA directly on the text embedding table. Directly mapping the text embedding table to the ID-based model’s output space, with only a basic projection head, leads to substantial semantic information loss.
  • Replacing the dot product with cosine in ROMA results in reduced performance in AppGallery dataset. So we leverage dot product as the similarity function, and it’s widely used in recommendation system, such as SASRec (Kang and McAuley, 2018), UniSRec (Hou et al., 2022).
  • The performance of MF2 w/o text is comparable to that of DIN, but significantly lags behind CELA. This indicates that CELA’s performance improvement benefits from semantic information.
These observations underscore the importance of both the DAP and ROMA stages in CELA for maximizing the PLM’s potential in recommendation systems, highlighting the critical role of adapting the PLM to domain-specific knowledge and tailoring it for task-specific objectives.
Table 4. Ablation study. The language model is RoBERTa and backbone is DIN.

### v3

@@ 4.2.1. ID-Based Model
4.2.2. Item-Level Modal Alignment
First, the textual and non-textual features on the item side are encoded into their respective representations. For item $i$, its non-textual features $\boldsymbol{x}^{i,n}$, are encoded utilizing the frozen ID embedding table developed in Section 4.2.1. Conversely, the textual feature $\boldsymbol{x}^{i,t}$ is processed through the PLM, followed by a transformation via a projection head. This head is to adjust the dimensionality of the output from the PLM to match that of the non-textual features. This encoding process is formulated as:
@@ 4.2.2. Item-Level Modal Alignment
where $\boldsymbol{e}^{n}_{i}$ represents the encoded representation of item $i$’s non-textual features, and $\boldsymbol{e}^{t}_{i}$ represents the encoded textual feature. The terms $\mathbf{W}$ and $\boldsymbol{b}$ are the parameters of the projection head, representing the matrix and bias, respectively.
Contrastive learning is then utilized to align the representations of textual and non-textual features on the item side. Within a given batch, representations of these features from the same item are considered positive pairs, while unmatched representations are treated as negative pairs. Given that both modalities depict the same item, there is an inherent overlap of insights between its positive pair, analogous to the relationship among visual, textual, and auditory data (Girdhar et al., 2023). By incentivizing the model to differentiate between matching and mismatched pairs, contrastive learning facilitates the establishment of correspondence or mapping functions across modalities. The alignment begins with
(10) $\small\mathcal{L}_{t,n}=-\frac{1}{\left|\mathcal{I}\right|}\sum_{i\in\mathcal{I}}\log\frac{\exp\left(sim\left(\boldsymbol{e}^{t}_{i},\boldsymbol{e}^{n}_{i}\right)/\tau_{2}\right)}{\sum_{k\in\mathcal{I}}\exp\left(sim\left(\boldsymbol{e}^{t}_{i},\boldsymbol{e}^{n}_{k}\right)/\tau_{2}\right)},$
where $\tau_{2}$ denotes another temperature scaling the similarity.
@@
(11) $\small\mathcal{L}_{n,t}=-\frac{1}{\left|\mathcal{I}\right|}\sum_{i\in\mathcal{I}}\log\frac{\exp\left(sim\left(\boldsymbol{e}^{n}_{i},\boldsymbol{e}^{t}_{i}\right)/\tau_{2}\right)}{\sum_{k\in\mathcal{I}}\exp\left(sim\left(\boldsymbol{e}^{n}_{i},\boldsymbol{e}^{t}_{k}\right)/\tau_{2}\right)}.$
The overall modal alignment loss combines these two components:
@@
Training is exclusively confined to the PLM and projection head, with the ID embedding table remaining non-trainable to prevent disruption of the collaborative filtering space. It is important to emphasize that, the item-level alignment significantly reduces the computational resources and time required for training.
@@ 4.3. MF2: Multi-Modal Feature Fusion
To mitigate suboptimal performance caused by purely text-based models, we preserve collaborative signals captured by the ID-based model while integrating the aligned PLM. Serving solely as an item textual feature encoder, the PLM avoids scalability issues from increased features and prevents information loss by encoding textual features of user behaviors item by item.
To reduce training time and online latency, we create an item text embedding table. Text for each item is processed through the aligned PLM to obtain the embedding, and stored for quick access by item identifiers, eliminating the need for real-time PLM processing. Then, as depicted on the right side of Figure 2, the recommender system is enhanced by integrating the item’s textual feature alongside non-textual features. This item textual feature, indexed from the text embedding table using the item’s identifier, is processed using an MLP to ensure dimensional consistency. Within the feature interactions layer, textual and non-textual features are treated uniformly. Such uniformity across all modalities prevents substantial modifications to the original network, thus facilitating its application to any network architecture. This process is formulated as:
@@
where $\mathbf{E}^{i,t}$ denotes the text embedding table, $\boldsymbol{x}^{i,id}$ denotes the candidate item’s ID, and $\boldsymbol{x}^{h,id}$ denotes items’ IDs in user history. $\boldsymbol{e}^{\prime}$ is an augmented version of $\boldsymbol{e}$, concatenated with textual embeddings of the candidate item and user history. $FI(\cdot)$ denotes the feature interaction layer and $FC(\cdot)$ denotes the fully connected layer.
@@ 4.4. Alternate Training
The last two stages, as indicated by the green dotted line in Figure 2, are designed to be executed alternately. Following the completion of the ID embedding table training in the MF2 stage, we can revert to the ROMA stage. Here, the item text representation is realigned to the newly trained ID embedding table. Subsequently, this realigned PLM is utilized to generate a new text embedding table. This updated table is then employed in the subsequent iteration of the MF2 stage for further training. This alternate approach allows for iterative refinement of both the alignment and fusion processes, thereby progressively improving the overall performance.
@@ 5.1.1. Datasets
We conduct offline experiments on three datasets, with details provided in Appendix 6.
@@ 5.1.2. Baselines
Our method is benchmarked against contemporary state-of-the-art models, encompassing both ID-based and text-enhanced models. Classic ID-based models, such as AutoInt (Song et al., 2019), DCN (Wang et al., 2017), DCNv2 (Wang et al., 2021), FiBiNET (Huang et al., 2019), PNN (Qu et al., 2016) , WideDeep (Cheng et al., 2016), DeepFM (Guo et al., 2017), and xDeepFM (Lian et al., 2018), focus on modeling high-order feature interactions. In contrast, models like DIN (Zhou et al., 2018) and DIEN (Zhou et al., 2019) specialize in fine-grained user behavior modeling. Text-enhanced models, including P5 (Geng et al., 2022) and PTab (Liu et al., 2022), aim to describe data in natural language to elicit responses from language models. CTR-BERT (Muhamed et al., 2021) integrates traditional and textual features for fusion in an MLP, UniSRec (Hou et al., 2022) utilizes a frozen BERT (Devlin et al., 2019) as its text encoder and CTRL (Li et al., 2023) aligns textual and tabular data representations to infuse the ID-based model with semantics. BoW (Zhang et al., 2010) uses bag-of-words representations as item text representations.
@@ 5.1.3. Evaluation Metrics
In this research, Area Under the Curve (AUC) and logarithmic loss (Logloss) are used as evaluation metrics for CTR models (Song et al., 2019; Wang et al., 2021; Qu et al., 2016; Guo et al., 2017). AUC assesses ranking ability, while Logloss measures predictive accuracy.
@@ 5.2. Overall Performance (RQ1)
Table 1. Performance comparison with baselines. The best results are in bold, second-best are in underlined. denotes statistical significance (p ¡ 0.05) compared to the second-best.
@@
  • PLMs enhance ID-based models as item textual feature encoders. UniSRec and CELA surpass ID-based models on MovieLens and AppGallery. This demonstrates that PLMs capture valuable information from item textual features not exploited by collaborative filtering, thereby improving model efficacy.
  • Purely text-based models typically underperform ID-based models. This is evidenced by the performance of P5, and PTab, which are inferior to most ID-based models. These findings are consistent with the results reported by (Li et al., 2023), which underscore the limitations of exclusive reliance on semantic modeling.
  • The integration of semantic and collaborative signals leads to improved performance. This is demonstrated by CTRL, UniSrec, and CELA, which excel in most cases. The injection of semantics into the ID-based model distinguishes them from traditional ID-based models. Furthermore, the retention of collaborative signals makes them outperform other text-enhanced models.
  • Task-oriented alignment is crucial for PLMs to eliminate the representation discrepancy and better adapt to downstream tasks. By aligning the outputs of PLMs and ID-based models, CTRL and CELA outperform UniSRec across three datasets. This outcome indicates that a simple whitening process in UniSRec is insufficient to bridge the gap. Therefore, alignment is essential to tailor PLMs for recommendation purposes.
  • Using PLMs solely as item textual feature encoders prevents information loss from feature proliferation. Compared to CTRL, CELA performs better in all cases since it extracts textual information from behavior sequences by encoding each item’s textual feature individually, thus avoiding lengthy annotations resulting from including all features of an item. In contrast, CTRL transforms tabular data into textual data using a hard prompt template and takes the PLM as a textual data encoder, which fails to provide detailed annotations for items in user behaviors due to token number limitations, leading to information loss.
@@ 5.3. Efficiency Measurement (RQ2)
Table 2. Comparison of training and inference efficiency between different models on AppGallery. Params is the number of trainable parameters, the inference time is calculated for one batch and the total training time is for the whole process.
Model Params (M) Inference (ms) Training (h)
@@
BoW 1.07 70.48 1.59
CELA DAP 102.88 / 1.07
CELA ROMA 102.88 / 2.06
@@
CELA / 69.64 5.42
To explore the practicality of CELA in industrial scenarios, we measure its training duration and inference latency on the industrial dataset AppGallery. Results presented in Table 2 reveals that:
  • Models utilizing hard prompt templates are impractical for real-world applications due to high training costs or inference overhead. This is attributable to their model complexity, indicated by the number of parameters. Training models like P5, PTab, CTR-BERT, and CTRL on extensive interaction-level textual data requires significant time, leading to infrequent updates and the use of outdated models, which negatively impacts performance. Additionally, the high inference overhead of P5, PTab, and CTR-BERT renders them unsuitable for real-time serving.
  • UniSRec and CELA maintain training and inference costs comparable to DIN. However, UniSRec’s performance is limited due to the lack of fine-tuning of the PLM. This limitation arises from insufficient alignment of the PLM’s knowledge with downstream tasks, resulting in suboptimal integration of textual and non-textual features. In contrast, CELA fine-tunes the PLM with negligible training expense, achieving superior performance.
  • CELA is efficient in both training and inference, making it industry-friendly. Its low training expense results from fine-tuning the PLM at the item level rather than the interaction level, as the number of items is significantly lower than the number of interactions, with a difference of up to 550 times here. Moreover, CELA’s inference cost is comparable to that of DIN as it avoids real-time text processing, utilizing an additional text embedding table that occupies only 87.2MB of storage space.
@@
To assess CELA’s efficacy in cold-start scenarios, the test dataset is segmented into groups based on item popularity, with the AUC evaluated for each group. As illustrated in Figure 3, CELA consistently outperforms the baseline across all cold-start conditions. This superior performance is attributed to the incorporation of a PLM that leverages embedded external knowledge for text comprehension. Notably, when the popularity falls below 20, the improvements of CELA become particularly pronounced. This is because, under such conditions, ID-based models struggle to learn effective ID embeddings owing to their dependence on historical interactions.
@@ 5.5. Ablation Study (RQ4)
To explore the impact of different training processes and components on the overall performance of CELA, we conduct an ablation study and the results are presented in Table 3. Firstly, we propose a variant to verify that the performance gains of CELA are attributed to its semantics rather than additional parameters:
  • Scaled DIN: Replacing CELA’s non-trainable text embedding table with a trainable one.
5.5.1. Domain-Adpative Pre-training
Domain-Adpative Pre-training (DAP) improves performance by enhancing the PLM’s understanding and expressiveness of dataset-specific texts. We design three variants to verify its effectiveness:
  • DAP w/o MLM: The MLM loss is removed in DAP.
  • DAP w/o SimCSE: The SimCSE loss is removed in DAP.
@@ 5.5.1. Domain-Adpative Pre-training
As shown in Table 3, removing either MLM or SimCSE losses, or both, results in performance degradation compared to the full CELA, demonstrating that both losses contribute to performance improvement. However, using only the SimCSE loss in pre-training, without the MLM loss, performs worse than no pre-training on the Amazon dataset. This indicates that SimCSE loss is more suitable as an auxiliary objective to MLM loss, as originally proposed in (Gao et al., 2021).
5.5.2. Recommendation-Oriented Modal Alignment
Recommendation-Oriented Modal Alignment (ROMA) boosts performance by aligning the outputs of the PLM with the ID-based model at the item level. We designed three variants to verify its effectiveness and explore what accounts for the enhancement:
  • CELA w/o ROMA: No alignment is applied, and the text embedding table in MF2 is encoded by a non-aligned PLM.
  • CELA w/o semantics: The text embedding table in MF2 is replaced by ROMA’s ID embedding table, excluding any semantics.
  • ROMA w/o PLM: The PLM in ROMA is replaced by a pre-encoded text embedding table, with only the table’s weights being directly updated.
Removing ROMA results in performance degradation, affirming the indispensability of task-oriented alignment of PLM outputs. The performance of CELA w/o semantics is comparable to that of DIN but significantly lags behind full CELA, indicating that CELA’s performance gains stem from the integration of semantics and imply negligible semantic loss during training. Moreover, directly applying ROMA to the text embedding table, instead of the original PLM, leads to reduced performance due to less preservation of semantics, evident from the reduced number of parameters. This underscores the necessity of PLM in ROMA.
Table 3. Ablation study. The language model is RoBERTa and the backbone is DIN.